OpenStudy (anonymous):

An investor invested a total of 2,100 in two mutual funds. One fund earned a 8%profit while the other earned a 5% profit. if the investors total profit was 129 how muchwas invested in the mutual fund? The amount investead in the mutual fund earned 8% was= ? The amount investead in the mutual fund earned 5% was=?

OpenStudy (anonymous):

@agent0smith can you please help me?

OpenStudy (whpalmer4):

Let x be the amount invested in the 8% fund. (2100-x) will be invested in the 5% fund. x*8% is the return on the 8% fund. (2100-x)*5% is the return on the 5% fund. The sum of the two returns is 129. Converting the percentages to decimals, we have \[x*0.08 + (2100-x)*0.05 = 129\] Solve that for \(x\) to find the amount invested in the fund that earned 8%. Subtract \(x\) from 2100 to find the amount in the fund that earned 5%. You might multiply both sides of the equation by 100 as a first step to get rid of the decimals.

OpenStudy (anonymous):

would x=2,075.8 @whpalmer4

OpenStudy (whpalmer4):

\[8x + 5(2100-x) = 12900\] \[8x -5x + 10500 = 12900\] \[3x = 2400\] x=?

OpenStudy (anonymous):

x=800 @whpalmer4 were did you get 12900 from

OpenStudy (anonymous):

nvm i saw my mistake i forgot to multiply 100 to 129 @whpalmer4

OpenStudy (whpalmer4):

when I multiplied both sides by 100

OpenStudy (whpalmer4):

How much is invested in the 5% fund?

OpenStudy (anonymous):

is that the 8% answer? @whpalmer4

OpenStudy (whpalmer4):

"Let x be the amount invested in the 8% fund. (2100-x) will be invested in the 5% fund."

OpenStudy (anonymous):

1,300 @whpalmer4

OpenStudy (whpalmer4):

Okay, 800 @ 8%, 1300 @ 5%

OpenStudy (anonymous):

Thank you so much! those are the answers right? @whpalmer4

OpenStudy (whpalmer4):

800 * 8% = 64 1300 * 5% = 65 64 + 65 = 129 Looks like they are correct.

OpenStudy (anonymous):

thank you so so so much! @whpalmer4 time for the next question hehe